A cannon launches a cannonball with mass 5.0 kg horizontally off a cliff that is 25.5 meter high. If the ball lands 64.537m from the edge of the cliff, find the muzzle velocity and momentum of the cannon shot. And if the cannon ball hits an enemy ship at close range(so you can assume the velocity is stilll the muzzle velocty) and penetrates it's hull stopping in 1/50 second, what was the force it hit with and find the average deceleration(negative of acceleration) during that half second and how deep it penetrates the hull.
Answer #1

time of flight

h = 0.5 gt^2

25.5 = 4.9 t^2

t = 2.281 s

motion along horizontal

64.537 = v * 2.281

v = 28.29 m/s

momentum = mv = 141.45 kgm/s

======

deceleration

a = v/ t= 28.29 / (1/50)

a = 1414.5 m/s^2

force = ma = 5* 1414.5 = 7072.5 N

======

Comment in case any doubt, will reply for sure.. Goodluck